Please help me solve this:
2b+a=11
a-b=5

Answer by xcentaur(357)

a+2b=11
a-b=5


multiply the second equation by 2.when you multiply an equation by a number,the equation still remains the same,because you could divide that and get the original equation back again


now we have
a+2b=11
2a-2b=10


Add the two equations,
(a+2b+2a-2b)=11+10
3a=21
a=7


From equation 1,
a+2b=11
2b=11-a
2b=11-7
2b=4
b=2


a=7,b=2


-Prabhat